OpenStudy (mbma526):

Can you please help me with this. Write an equation of a line that passes through the point (0, 5).

OpenStudy (mbma526):

@TheAnarchist

OpenStudy (anonymous):

y=x+5

OpenStudy (theanarchist):

Y=(9/7)*X+5 Hope this helps

OpenStudy (mbma526):

I think it is y= 9 over 7 . x +5

OpenStudy (mbma526):

Thats what I got and it does @TheAnarchist

OpenStudy (theanarchist):

yeah that's what I said (9/7) is the same thing as Y= 9 over 7, just sayin

OpenStudy (mbma526):

Yeah @TheAnarchist and thank you